OpenStudy (anonymous):

Find the hypotenuse of each isosceles right triangle when the legs are of the given measure. Given = 6√2

OpenStudy (anonymous):

= how many inches?

OpenStudy (anonymous):

See: In case of Isosceles right triangle: \[a^2 + a^2 = H^2 \implies H = \sqrt{2a^2} \implies H = a \sqrt{2}\] Here a = \(6\sqrt{2}\). Can you find H here??